# Stratified Selective Sampling for Instruction Tuning with Dedicated Scoring Strategy

### v1

@@ 3.2 Scoring
For each $(x_{i},y_{i})\in\mathcal{D}$, we compute the preference score $p_{i}=f_{i}\cdot q_{i}$, where $f_{i}$ and $q_{i}$ are the difficulty and quality scores, respectively. Each of them is normalized using min-max scaling, with the 1st and 99th percentiles as the minimum and maximum values across all samples in $\mathcal{D}$. The quality scores are normalized per scorer ($q_{math}$, $q_{code}$, $q_{if}$ and $q_{deita}$) as they have differing ranges. For multi-turn conversations, where each data point $d_{i}$ consists of a sequence of turns $\{(x_{0},y_{0}),...,(x_{T},y_{T})\}$, we assign a category $l_{t}$ to each turn and determine the conversation-level category $l_{i}$ by either selecting the most frequent category or defaulting to the first one. We then compute turn-level difficulty and quality scores, $f_{t}$ and $q_{t}$, based on their respective categories $l_{t}$. These are averaged across all turns to yield the overall conversation-level scores $f_{i}$ and $q_{i}$. [^3] For a given conversation, quality scores are averaged separately for each category; $q_{i}$ is then selected according to the main category $l_{i}$.
3.3 Sampling
Greedily choosing the highest-scoring samples often leads to redundancy in some domains and under-representation in others. Thus, maintaining diversity in the final dataset $\mathcal{D}^{\prime}$ is essential. Since diversity is a property of the dataset as a whole—not of individual samples—selection should consider the dataset globally (e.g., Ge et al., 2024), rather than relying on iterative, sample-by-sample strategies (e.g., Liu et al., 2024b; Bukharin et al., 2024). Diversity can be promoted top-down by balancing category proportions (see e.g. Grattafiori et al., 2024b; Dong et al., 2024), or bottom-up by ensuring sufficient semantic dissimilarity among selected samples (e.g., Liu et al., 2024b; Ge et al., 2024; Lu et al., 2024).
@@ 4.1 Experimental Setup
We use the IT datasets detailed in Table 2. In our experiments, the full IT dataset $D$ is the aggregation of all listed datasets.
Dataset #samples (#turns) HuggingFaceH4/ifeval-like-data 5K vicgalle/alpaca-gpt4 52K nvidia/OpenMathInstruct-2555Without augmented problems. 52K ai2-adapt-dev/flan_v2_converted 90K openbmb/UltraInteract_sft (Coding) 115K WizardLMTeam/WizardLM_evol_instruct_V2_196K 143K theblackcat102/sharegpt-english 50K (392K) microsoft/orca-agentinstruct-1M-v1666Randomly sampled. 200K (903K) all 707K (1.75M)
Table 2: Instruction tuning dataset overview.
Figure 2: Category proportions of different sampling strategies for 100k samples.
@@
While most experiment results use finetuned Mistral-7B-v0.3, we also demonstrate generalization across models of varying sizes and families: i) tiiuae/Falcon3-10B-Base, ii) meta-llama/Llama-3.1-8B, iii) mistralai/Mistral-7B-v0.3, iv) Qwen/Qwen2.5-3B and v) HuggingFaceTB/SmolLM2-1.7B.
(a) Aggregation models & benchmarks
(b) Aggregation benchmarks
(c) Aggregation models
Figure 3: Performance gains over the base model for different sampling strategies with 100k samples. We aggregate the results (a) across all 5 tested models and all 13 benchmarks; (b) across all benchmarks separated by models; (c) across all models separated by benchmarks.
@@
We compare against the following baseline methods for constructing $D^{\prime}\subset D$: (i) Random, where we sample uniformly at random, (ii) Longest (Zhao et al., 2024a), where we include samples having the longest responses and (iii) Deita (Liu et al., 2024b), which ranks data points according to complexity and quality scores ($o_{deita}$ and $q_{deita}$, resp.), and iteratively builds $D^{\prime}$ by adding samples that are dissimilar to those already selected.
@@
We evaluate the success of our instruction tuning experiments by testing Models on a suite of benchmarks, covering a broad spectrum of model capabilities and giving a holistic picture of the tuning success. All tested benchmarks are listed in Table 11 in Appendix A.7.
4.2 Main Results
For every selection strategy we construct a subset $\mathcal{D}^{\prime}={d}_{i=1}^{m}\subset\mathcal{D}$ with size $m=100,000$ and subsequently fine-tune each of the five base models on $\mathcal{D}^{\prime}$. In addition to the baseline strategies described in Section 4.1, we investigate four principled variants that exploit the pipeline components introduced in Section 3.2:
  1. i)
    Quality—sort $\mathcal{D}$ by $q_{i}$ and select the top $m$ items;
  2. ii)
    Difficulty—sort by $f_{i}$ and select the top $m$ items;
  3. iii)
    Combination—sort by the preference score $p_{i}=f_{i}\cdot q_{i}$ and select the top $m$ items;
  4. iv)
    Combination++—as in combination, but with sampling via the clustering-and-quota procedure from Section 3.3.
Figure 2 shows task-type distributions across sampling strategies, including baselines. Figure 3(a) reports average performance gains over the untuned base model for all evaluation benchmarks. The full pipeline (combination++) achieves the highest overall performance, surpassing every baseline and even the model trained on the entire source set with $|\mathcal{D}|>707\text{k}$.
To examine robustness, we break the results down by model size (Figure 3(b)). Combination++ yields the most consistent gains of all tested conditions for all five bases and is outperformed by full-data training in only two cases, confirming that the proposed sampling approach transfers well across models. A benchmark-wise analysis (Figure 3(c)) shows similarly stable improvements, with combination++ delivering consistent performance throughout.
(a) MT-Bench
(b) Alpaca Eval 2.0
Figure 4: Performance on LLM-as-a-judge benchmarks.
For completeness, we also evaluate our approach on benchmarks using LLM-as-a-judge. Due to the high cost of the proprietary LLM-judges, we limit ourselves to evaluating models trained on Mistral-7B-Base under a single experimental condition. As shown in Figure 4, for $m=100,000$, combination++ performs best in the length-controlled AlpacaEval 2.0 (Dubois et al., 2024), and matches random sampling on MT-bench (Zheng et al., 2023b).
Figure 5: Average benchmark score gains over the base model, when finetuning Mistral-7B-Base with different dataset sizes sampling using various strategies.
Robust downscaling
@@ 4.2 Main Results
We next test whether the full pipeline’s benefits persist when the target size $m$ is varied. Fixing the base model to Mistral-7B-Base, we sample additional datasets $\mathcal{D}^{\prime}$ of 1k, 10k, 25k, 50k items with combination++ and all baselines. Figure 5 demonstrates that our method outperforms the alternatives across every scale: notably, with only $m=50,000$ (7% of the source data), it surpasses the performance resulting from full-data training. For benchmark-specific plots, see Appendix A.8.1.
Figure 6: (a) Category proportions of $\mathcal{D}_{\text{skewed}}$ and the 25k subsets sampled with various selection strategies; (b) Average benchmark scores, when finetuning Mistral-7B-Base with data sampled from the skewed distribution using various strategies.
We, again, evaluate on LLM-as-a-judge benchmarks under limited setups. As before, combination++ performs comparably to other strategies on MT-bench, while outperforming them on AlpacaEval for $m=5,000$ and on average performance for $m=25,000$.
The surprising insensitivtiy of MT-Bench to sampling strategies may stem from the consistent presence of GPT-4-generated responses in every sampled datasets (see Figure 10, Appendix A.5), which is known to strongly influence MT-Bench outcomes (Panickssery et al., 2024; Stureborg et al., 2024; Wataoka et al., 2024).
Robustness to skewed data distributions
@@
We now probe how selection strategies perform when the candidate pool $\mathcal{D}$ is itself highly imbalanced. To simulate this, we create a skewed source set $\mathcal{D}_{\text{skewed}}\subset\mathcal{D}$ by randomly selecting two categories $l_{1},l_{2}\in\mathcal{L}$ and retaining only items with labels $l_{i}\in{l_{1},l_{2}}$, along with a small (4%) residue from all other categories. This yields a strongly biased data distribution of size $N_{\text{skewed}}=366\text{k}$ (see Figure 6(a) bottom, for the resulting category distribution).
From $\mathcal{D}_{\text{skewed}}$, we sample $m=25\text{k}$ items with the same strategies as previously and fine-tune Mistral-7B-Base on each $\mathcal{D}_{\text{skewed}}^{\prime}$. We expected diversity-aware sampling strategies like Deita and combination++ to mitigate bias by enforcing either semantic spread and explicit quotas. However, none of the strategies show a clear advantage, with random sampling performing with a slight, insignificant edge. On closer inspection of the category distributions of the sampled data in Figure 6(a), we find that sampling strategies without explicitly encouraging diversity, such as quality and difficulty, mostly maintain the bias of $\mathcal{D}_{\text{skewed}}$, while Deita exacerbates it. Combination++ successfully debiases the sample; however, this balancing does not translate into improved model performance.
@@ 5 Discussion and Conclusion
In this paper, we address this challenge by explicitly covering the whole spectrum of possible instruction domains and tailoring scoring strategies that are apt to judge the utility of samples in those domains. We show the robustness of our approach by testing it across diverse settings (model families, scales, and sample sizes) and evaluating it on a wide range of common benchmarks. Our experiments provide further evidence for the necessity of data selection, as our sampling not only significantly reduces the amount of data required for training, but also outperforms setups trained on the full source data. Despite its more elaborate design—including multiple scoring methods—our pipeline remains computationally efficient due to targeted scoring.
While our pipeline outperforms baselines in almost all settings, it shows no significant gains when applied to a strongly skewed source distribution. We caution against over-interpreting this outcome, as we only evaluate on a single type of distributional bias (toward two of seven categories). A more complete robustness assessment would require testing across various bias configurations, which is computationally infeasible and thus left for future work.
@@
A major issue in collecting data for our difficulty scorer is a certain unreliability in the evaluation of model responses. Previous research shows that evaluations oftentimes have weak robustness to e.g. the prompt formatting (Sclar et al., 2024; Weber et al., 2023a, b; Polo et al., 2024), bias of LLMs-as-a-judge (Panickssery et al., 2024; Stureborg et al., 2024; Wataoka et al., 2024) or errors in post-processing (such as issues in extracting the answer from the model response). For example, GPT4o showed overall weaker performance on some of our evaluated subsets than some small open-source models. Upon closer inspection, we encountered that – while providing the correct answer – GPT4o generally tends not to follow the formatting of the given few-shot examples and rather responds in an open-form manner, resulting in failing the tight response search masks of the used evaluation frameworks. While we try to mitigate this issue as much as possible, we cannot guarantee that difficulty scores exactly reflect a model’s capacity to solve a given data point.
@@
We did not develop nor have dedicated quality scorers for samples belonging to Reasoning, Factual QA and Extraction categories. However, we hypothesis that process reward models (PRMs) could be adapted to Reasoning tasks beyond math, such as spatial (e.g., Wu et al., 2024) and deductive reasoning (e.g., Seals and Shalin, 2024), given appropriate training data. Factuality assessment is a long-standing research problem that has become more relevant in the era of LLMs. Wei et al. (2024) introduced SAFE (Search-Augmented Factuality Evaluator), an LLM-agent-based method for automatically assessing long-form factuality in model response. Although significantly cheaper than human annotators (up to 20×20$\times$20 ×), SAFE still incurs costs of \$20–\$40 per 100 prompt-response pairs. For Extraction tasks, future work might draw inspiration from recent advances in RAG evaluation (Yu et al., 2025). These directions, however, are beyond the scope of this paper.
Robustness to skewed data.
Our results indicate that we cannot get performance gain in setups with very skewed data. However, the experimental setup is very limited, testing only skew in a single variation and more experiments are needed to draw conclusions. Further, there are many other skews that can occur in practice that we did not evaluate our sampling pipeline against.

### v2

@@ 3.3 Sampling
For each $(x_{i},y_{i})\in\mathcal{D}$, we compute the preference score $p_{i}=f_{i}\cdot q_{i}$, where $f_{i}$ and $q_{i}$ are the difficulty and quality scores, respectively. $f_{i}$ and $q_{i}$ are normalized per scorer as they have differing ranges, using min-max scaling, with the 1st and 99th percentiles as the minimum and maximum values across all samples in $\mathcal{D}$. For multi-turn conversations, where each data point $d_{i}$ consists of a sequence of turns $\{(x_{0},y_{0}),...,(x_{T},y_{T})\}$, these scores are averaged across all turns to yield the overall conversation-level scores $f_{i}$ and $q_{i}$.
Clustering + Ranking.
Greedily choosing the highest-scoring samples often leads to redundancy in some domains and under-representation in others. Thus, maintaining diversity in the final dataset $\mathcal{D}^{\prime}$ is essential. Since diversity is a property of the dataset as a whole—not of individual samples—selection should consider the dataset globally (e.g., Ge et al., 2024), rather than relying on iterative, sample-by-sample strategies (e.g., Liu et al., 2024c; Bukharin et al., 2024). Diversity can be promoted top-down by balancing category proportions (see e.g. Grattafiori et al., 2024; Dong et al., 2024), or bottom-up by ensuring sufficient semantic dissimilarity among selected samples (e.g., Liu et al., 2024c; Ge et al., 2024; Lu et al., 2024).
@@
Given the large value of $m_{l}$ and the high-dimensional nature of the input embeddings, we employ MiniBatchKMeans from scikit-learn with the k-means++ initialization method and a batch size of 2048.
Sampling evaluation.
We evaluate the effectiveness of our stratified sampling and clustering as we did for the difficulty and quality scorer evaluation in Sections 3.2.1 and 3.2.2. In addition, we demonstrate the efficiency and robustness of Laser’s sampling technique, particularly in comparison with Deita (Liu et al., 2024c) and CaR (Ge et al., 2024). The respective results are resported in Appendix A.4.1.
@@ 4.1 Experimental Setup
We use the aggregation of all the listed datasets in Table 2 as the source dataset $\mathcal{D}$ in all experiments, if not specified otherwise.
@@
While most experiments we finetune Mistral-7B-v0.3, we also demonstrate generalisation across models of varying sizes and families: i) tiiuae/Falcon3-10B-Base, ii) meta-llama/Llama-3.1-8B, iii) mistralai/Mistral-7B-v0.3, iv) Qwen/Qwen2.5-3B and v) HuggingFaceTB/SmolLM2-1.7B.
(a) Score gains per benchmark category
(b) Score gains per model size
Figure 3: Performance gains over the base model for different sampling strategies with 100k samples. We aggregate the results (a) across all benchmarks separated by models; (b) across all models separated by benchmark categories.
@@
We compare against the following alternative methods for constructing $D^{\prime}\subset D$:
  1. i)
    random—sampling uniformly at random;
  2. ii)
    longest—including samples having the longest responses; [^6] In the case of multi-turn samples, we only consider the last turn’s responses.
  3. iii)
    deita (Liu et al., 2024c)—ranking data points according to complexity and quality scores ($o_{deita}$ and $q_{deita}$, resp.), and iteratively builds $D^{\prime}$ by adding dissimilar samples based on embedding distances.
[…]
Table 3: Evaluation benchmarks with associated category and evaluation details.
@@
We evaluate the instruction-tuned models on a suite of benchmarks listed in Table 3, leveraging popular evaluation frameworks LM-evaluation harness (Gao et al., 2024) and AlpacaEval [^7] We use openai/gpt-oss-120b as the LLM-judge. (Dubois et al., 2024). We report normalized score gain as the main evaluation metric, computed as the performance improvement over base models, scaled with min-max normalization.
4.2 Results
4.2.1 Main results
For each selection strategy, we construct a subset $\mathcal{D}^{\prime}={d}_{i=1}^{m}\subset\mathcal{D}$ of size $m=100,000$, and subsequently fine‑tune each of the five base models on $\mathcal{D}^{\prime}$. Figure 2 shows the resulting category distributions across sampling strategies, including the baselines. Figure 3(a) reports normalized score gains over the untuned base models for all evaluation benchmarks and models. Laser achieves the highest overall performance, surpassing all baselines and even the models trained on the entire source set ($|\mathcal{D}|>707\text{k}$). We examine the robustness by breaking the aggregated results down by model size (Figure 3(b)). Laser yields the most consistent gains of all tested conditions for all five bases and outperforms the full‑data training in all but the smallest model, confirming that the proposed sampling approach transfers well across models.
Figure 4: Average score gains over the base model when ablating Laser with 100k samples: with only $f$, $q$ and $p$ as ranking criteria.
To better understand the contribution of each Laser component, we ablate sampling by ranking with only difficulty scores ($f$), quality scores ($q$), or their combination (pp)—without clustering or stratification—and compare these variants to the full pipeline. As shown in Figure 4, using $f$ or $q$ alone yields clear improvements over random sampling. However, their full potential is realized only when combined within the complete Laser pipeline. Notably, relying solely on $p$ for ranking provides only marginal gains, and even underperforms $f$- or $q$-based sampling individually, highlighting the critical role of stratified sampling and clustering in Laser.
Hence, all three properties (difficulty, quality and diversity) are necessary but not sufficient on their own for effective and robust data sampling. This is intuitive: highly challenging queries answered incorrectly provide little benefit, while accurate responses to trivial questions yield weak training signals. The insufficiency of using solely $f$ or $q$ is especially interesting when we consider their covariance: in our source data $\mathcal{D}$, quality and difficulty are generally negatively correlated, with high negative correlations in Coding (r=.38r=-.38) and Math (r=.43r=-.43), and smaller correlations in other categories. This explains the large performance gaps observed for Coding and Math in Figure 4. We observe a similar relationship for the source datasets $\mathcal{D}_{weak}$ and $\mathcal{D}_{strong}$ in the following section.
4.2.2 Additional robustness experiments
We continue to demonstrate the robustness of the full Laser pipeline for different target sample sizes $m$ and variations in the source data distribution $\mathcal{D}$, such as differences in source data quality and skewed data distributions.
Downscaling $m$.
@@ 4.2.2 Additional robustness experiments
As different users may require instruction data sets of different sizes, a selection pipeline should improve sampling at every target scale. To verify this, we fix the base model to Mistral‑7B‑Base and constructs datasets $\mathcal{D}^{\prime}$ of {1k, 5k, 10k, 25k, 50k} items using Laser as well as all baselines. Figure 5 demonstrates that our method outperforms the alternatives across scales, except when $m=1,000$. Notably, with only $m=50,000$ (7% of the source data), it surpasses the performance of full-data fine-tuning. We provide benchmark-specific plots in Figure 24 (Appendix A.7), reported without normalizing score gain.
Figure 5: Average score gains over the base model, when finetuning Mistral-7B-Base using instruction datasets with varying sample sizes.
Effectiveness shifts in distribution of $\mathcal{D}$.
Next, it is very probable that the average effectiveness of the source data $\mathcal{D}$ changes from scenario to scenario. We therefore test Laser with two shifted source distributions: $\mathcal{D}_{\text{strong}}$ consisting of the full Tülu v3 (Lambert et al., 2024) – a meticulously optimised dataset, which can be considered highly effective – and $\mathcal{D}_{\text{weak}}$ composed of multiple datasets that have previously proved to be less effective for instruction tuning (see Appendix A.5.2 for details).
We sample $m\in\{1\text{k},5\text{k},10\text{k},25\text{k},50\text{k},100\text{k}\}$ data points from both sources $\mathcal{D}_{\text{strong}}$ and $\mathcal{D}_{\text{weak}}$ using Laser and random sampling. The results of finetuning Mistral-7B-v0.3 are shown in Figure 6. Generally, Laser performs robustly on different source datasets, with fewer fluctuations when changing the sampling size and comparably more improvements for the stronger than for the weaker source dataset. Interestingly, the performance gains using Laser over random are more pronounced for $\mathcal{D}_{\text{weak}}$ when sampling smaller target datasets, potentially distilling the few better examples from the weaker source.
Figure 6: Performance on source datasets of different effectiveness ($\mathcal{D}_{\text{weak}}$ and $\mathcal{D}_{\text{strong}}$) for different sample sizes. For comparison, “all” shows performance when using full $\mathcal{D}_{\text{weak}}$ or $\mathcal{D}_{\text{strong}}$ without sampling.
Skewed distribution of $\mathcal{D}$.
@@
Ultimately, $\mathcal{D}$ might be very homogeneous (i.e. contain mostly a specific data type). To simulate this, we create 5 skewed source sets $\mathcal{D}_{\text{skewed}}$ by randomly selecting two categories $l_{1},l_{2}\in\mathcal{L}$ for each $\mathcal{D}_{\text{skewed}}$ and retaining only items with labels $l_{i}\in{l_{1},l_{2}}$, along with a small (4%) residue from all other categories, yielding a strongly biased data distribution. We show the proportions of each category for all 5 new source datasets in Figure 7(a).
From each $\mathcal{D}_{\text{skewed}}$, we sample 25k items with the same strategies as previously and fine‑tune Mistral‑7B‑Base on each of the resulting $\mathcal{D}_{\text{skewed}}^{\prime}$. We expected diversity-aware sampling strategies like Deita and Laser to mitigate bias by enforcing either semantic spread and explicit quotas. Indeed, the diversity-aware sampling strategies on average outperform the naive methods, with Laser having an edge over Deita, in terms of performance improvements in Figure 7(b) Upon closer inspection, we find that only Laser is reducing the introduced bias of categories while Deita exacerbates it (see details in Appendix A.5.2). Next to that, we observe a key limitation of Deita’s sampling method in this setup: Only adding dissimilar examples leads to quick saturation when source data is homogeneous and, thereby, hitting a ceiling of possible samples that can be sampled with Deita.
Figure 7: (a) Category proportions of different $\mathcal{D}_{\text{skewed}}$ source distributions; (b) Average benchmark scores, when finetuning Mistral-7B-Base with data sampled from the skewed distribution using various strategies.
@@ 5 Discussion and Conclusion
In this paper, we address this challenge by explicitly covering the whole spectrum of possible instruction domains and tailoring scoring strategies that are apt to judge the utility of samples in those domains. We show the robustness of our approach by testing it across diverse settings (model families, scales, various shifts in the source data distribution) and evaluating it on a wide range of common benchmarks. Our experiments provide further evidence for the necessity of data selection, as our sampling not only significantly reduces the amount of data required for training, but also outperforms setups trained on the full source data. Despite its more elaborate design—including multiple scoring methods—our pipeline remains computationally efficient due to targeted scoring.
While our pipeline outperforms baselines in almost all settings, it shows only little improvements when sampling large portions from weak source data (as shown in Section 4.2.2). This might be due to a potential ceiling effect on possible performance given the weak source material, which might require improving the source data rather than elaborate sampling techniques. The modular setup of Laser allows for further incremental improvements of different scorers in future work.
@@
A major issue in collecting data for our difficulty scorer is a certain unreliability in the evaluation of model responses. Previous research shows that evaluations oftentimes have weak robustness to e.g. the prompt formatting (Weber et al., 2023a, b; Sclar et al., 2024; Polo et al., 2024b), bias of LLMs-as-a-judge (Panickssery et al., 2024; Stureborg et al., 2024; Wataoka et al., 2024) or errors in post-processing (such as issues in extracting the answer from the model response). For example, GPT4o showed overall weaker performance on some of our evaluated subsets than some small open-source models. Upon closer inspection, we encountered that – while providing the correct answer – GPT4o generally tends not to follow the formatting of the given few-shot examples and rather responds in an open-form manner, resulting in failing the tight response search masks of the used evaluation frameworks. While we try to mitigate this issue as much as possible, we cannot guarantee that difficulty scores exactly reflect a model’s capacity to solve a given data point.
@@
We did not develop nor have dedicated quality scorers for samples belonging to Reasoning, Factual QA and Extraction categories. However, we hypothesize that process reward models (PRMs) could be adapted to Reasoning tasks beyond math, such as spatial (e.g., Wu et al., 2024) and deductive reasoning (e.g., Seals and Shalin, 2024), given appropriate training data. Factuality assessment is a long-standing research problem that has become more relevant in the era of LLMs. Wei et al. (2024) introduced SAFE (Search-Augmented Factuality Evaluator), an LLM-agent-based method for automatically assessing long-form factuality in model response. Although significantly cheaper than human annotators (up to $20\times$), SAFE still incurs costs of \$20–\$40 per 100 prompt-response pairs. For Extraction tasks, future work might draw inspiration from recent advances in RAG evaluation (Yu et al., 2025). These directions, however, are beyond the scope of this paper.
Performance with weak source data.
Our results indicate that we get only small performance gains in setups with very weak source data. However, there might be a ceiling for possible performance in such settings, as data selection alone might not be sufficient to achieve good performance in such setups.
Acknowledgments
We want to thank Rishiraj Saha Roy, Lucas Druart and Christian Kroos for their valuable feedback on earlier versions of this manuscript. This work was partly funded by the German Federal Ministry for Economic Affairs and Climate Action (BMWK) through the project OpenGPT-X (project no. 68GX21007D) and partly by the Free State of Bavaria in the DSgenAI project (Grant Nr.: RMF-SG20-3410-2-18-4). Further, the authors gratefully acknowledge the Centre for Information Services and High Performance Computing [Zentrum für Informationsdienste und Hochleistungsrechnen (ZIH)] and ScaDS.AI at the Technical University of Dresden for providing its facilities for experimental work.